OpenStudy (anonymous):

Convert x = 4 to polar form.

OpenStudy (anonymous):

r=

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

\[ x = r\cos(\theta) \]

OpenStudy (anonymous):

\[ x=4=r\cos \theta\implies r = 4\sec \theta \]

OpenStudy (anonymous):

so r=4/sec(theta)?